OpenStudy (anonymous):

Diameter=8.5in find the radius adn area

OpenStudy (anonymous):

radius is 1/2 diam d = 8.5 r = 4.25 a = pi r^2 a = 3.14 (4.25)^2 plug into calulator if the answer includes symbol for PI just do not multiply the 3.14